# APT-Pipe: An Automatic Prompt-Tuning Tool for Social Computing Data Annotation

### 24 Jan 2024 v1

@@ Abstract.
Recent research has highlighted the potential of LLM applications, like ChatGPT, for performing label annotation on social computing text. However, it is already well known that performance hinges on the quality of the input prompts. To address this, there has been a flurry of research into prompt tuning — techniques and guidelines that attempt to improve the quality of prompts. Yet these largely rely on manual effort and prior knowledge of the dataset being annotated. To address this limitation, we propose APT-Pipe, an automated prompt-tuning pipeline. APT-Pipe aims to automatically tune prompts to enhance ChatGPT’s text classification performance on any given dataset. We implement APT-Pipe and test it across twelve distinct text classification datasets. We find that prompts tuned by APT-Pipe help ChatGPT achieve higher weighted F1-score on nine out of twelve experimented datasets, with an improvement of 7.01% on average. We further highlight APT-Pipe’s flexibility as a framework by showing how it can be extended to support additional tuning mechanisms.
Human computation, crowdsourcing, prompt-tuning, large language models, data annotation
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; May 13–17, 2024; Singapore
@@ 1. Introduction
Data annotation commonly relies on human intelligence and suffers from challenges related to annotator quality, data comprehension, domain knowledge, and time cost (Haq et al., 2022; Huang et al., 2016; Veit et al., 2016; Chalkidis et al., 2020). Recent research has highlighted that prompt-based LLM applications, like ChatGPT, have the potential to substitute human intelligence as a data annotator for social computing tasks, e.g., sentiment analysis (Bang et al., 2023; Zhu et al., 2023) and hate speech detection (Huang et al., 2023). Use of tools like ChatGPT for such annotation tasks can significantly improve the cost and speed for researchers. However, such models are heavily impacted by prompt quality, whereby poorly tuned prompt can severely damage annotation accuracy (Shum et al., 2023). This creates the need for tools that can help researchers design and evaluate effective prompts.
@@
To date, most prompt-tuning for annotation tasks is done either manually (Liao et al., 2022) or by injecting prior knowledge of the dataset (Gilardi et al., 2023; Huang et al., 2023). Prompt-tuning for ChatGPT, however, usually introduces demands on expertise that can be hard to address. Thus, we argue that an automated prompt-tuning tool for ChatGPT tasks could remove a significant manual load related to learning prior knowledge of the dataset to annotate. Researchers have already begun to propose techniques for such automation (Petroni et al., 2019; Min et al., 2021; Bang et al., 2023). We focus on three early-stage ideas: introducing standardized templates that (almost) guarantee consistent responses; introducing few-shot examples into prompts to guide annotations; and embedding additional text metadata to augment raw text information. In practice, however, implementing these simple concepts raises key challenges. First, ChatGPT can generate unpredictable and unstructured responses that make parsing annotations difficult (as its responses lack a standard format). Although cloze-style prompts can improve responses’ parsability, such templates usually introduce a trade-off on ChatGPT’s time efficiency (Petroni et al., 2019). Second, although preliminary evidence suggests that performance can be improved by including few-shot examples of annotation (Min et al., 2021), manual exemplar selection can be time consuming. Thus, techniques need to be developed to automatically identify suitable few-shot examples for the prompt. Third, although embedding natural language processing (NLP) metrics in prompts (Bang et al., 2023) has been shown to increase performance, the selection of such metrics requires domain specific expertise. Thus, techniques need to be developed to select the best permutations of NLP metrics to include in prompts. Importantly, the suitability of the above approaches may differ on a per-dataset basis, making a single standardized approach impractical.
@@
To address these challenges, we propose the Automatic Prompt-Tuning Pipeline (APT-Pipe), an extensible prompt-tuning framework for improving text-based annotation task performance. APT-Pipe only requires humans to annotate a small sample subset of data to offer a robust ground truth. It then exploits this annotated subset to automatically tune and test prompts with the goal of improving ChatGPT’s classification performance. APT-Pipe addresses the above three challenges using a three-step prompt-tuning pipeline. It first generates JSON template prompts that can make ChatGPT respond consistently in a desired format. APT-Pipe then automatically computes the most suitable few-shot exemplars to include in prompts. Finally, through a number of iterations, APT-Pipe tests a variety of prompt configurations with NLP metrics to identify which work best for the particular dataset. These three steps are derived from our experience and state-of-the-literature. However, we acknowledge that many other prompt-tuning techniques are likely to emerge in the coming years. Consequently, we design APT-Pipe in an extensible and modular fashion, allowing other researchers to easily “plug-in” additional tuning strategies.
@@
We implement APT-Pipe with the ChatGPT gpt-3.5-turbo model and evaluate it across twelve distinct text classification datasets drawn from three domains in social computing. APT-Pipe will be made publicly available for use. Our results shows that prompts tuned by APT-Pipe can improve ChatGPT’s classification performance on nine out of twelve datasets, with a 7.01% increase of weighted F1-score on average. Moreover, prompts tuned by APT-Pipe can decrease ChatGPT’s computing time for classification, while providing over 97% of all responses in a consistent and easy-to-parse format. Finally, we highlight APT-Pipe’s flexibility by extending its framework with two state-of-art tuning mechanisms called Chain-of-Thought and Tree-of-Thought.
The contributions of this paper can be summarized as follows:
@@
  • We show that APT-Pipe prompts generate annotations in a consistent format and take less time compared to existing prompt designs. On average, 97.08% responses generated by APT-Pipe are parsable by desired format and it decreases more than 25% time cost of the baselines.
  • We highlight APT-Pipe’s extensibility by introducing two additional prompt-tuning techniques, Chain of Thought (CoT) and Tree of Thought (ToT). We show that CoT improves APT-Pipe’s F1-scores on five datasets by 4.01%, and ToT improves APT-Pipe’s F1-scores on six datasets by 3.49%. Importantly, we show that APT-Pipe can learn these improvements on a per-dataset basis, automatically tuning the prompts without human intervention.
@@ 2.1. LLMs in Social Computing
Social computing research relies extensively on human-annotated datasets. Such annotations are time consuming and often costly. The release of LLM tools such as ChatGPT (Bang et al., 2023; Guo et al., 2023) has uncovered a range of possibilities to automate text data labeling tasks due to their robust performance in natural language comprehension and reasoning (Kojima et al., 2022).
Recently, some studies (Sallam et al., 2023; Huang et al., 2023) have employed LLMs to assist in text annotation in social computing. Sallam et al. (Sallam et al., 2023) employ LLMs to discriminate misinformation about vaccines on social media platforms. Huang et al. (Huang et al., 2023) exploit LLMs to annotate implicit hateful tweets and report nearly 80% accuracy. Zhu et al. (Zhu et al., 2023) evaluate the potential of LLMs to annotate five different types of text data. However, work has revealed that the annotation performance of LLMs varies across different LLMs and different domains and datasets (Aiyappa et al., 2023; Shen et al., 2023). The above literature demonstrates that, although it is feasible to employ LLMs for text annotation tasks, the generality and reliability of LLMs need to be improved.
@@ 2.2. Prompt-Tuning for Text Annotation
In existing LLM-based text annotation methods (Alizadeh et al., 2023; Luo et al., 2023; Kuzman et al., 2023; Reiss, 2023), several prompt-tuning techniques have been utilized to enhance the performance of LLMs. Alizadeh et al. (Alizadeh et al., 2023) investigate the performance difference between LLMs and crowd-workers in text annotation tasks. They employ natural language formatted prompts in zero-shot tests. For few-shot tests, they employ the Chain-of-Thought (CoT) (Wei et al., 2022) prompting technique, where LLMs are provided with the question and step-by-step reasoning answers as examples. Kuzman et al. (Kuzman et al., 2023) conduct genre identification tasks using LLMs by feeding prompts with manually defined task descriptions, and manually extracting the answers from responses. Reiss (Reiss, 2023) applies LLMs to text annotation tasks by manually designing 10 prompts, feeding them to LLMs, and evaluating their performance. From the above literature, we observe that existing prompt-tuning methods require the participation of domain expertise and human judgement, which hinders the utilization of LLMs in automated annotation tasks. Using ChatGPT as a case-study, we focus on proposing a pipeline that can automatically perform prompt tuning.
@@ 3. APT-Pipe Design
We now describe our pipeline for tuning a text classification prompt. The pipeline’s goal is to generate a tuned prompt to improve ChatGPT’s classification performance on a per-dataset basis. The pipeline operates in discrete steps, and we have designed our framework in an extensible fashion, allowing future modifications to be “plugged-in” as new steps. The current implementation has three steps. First, the prompt engineer must personalize a JSON-based prompt following our template (Step 1). Following this, the pipeline searches for suitable few-shot learning examples to include in the prompt (Step 2) and augmentary text metrics that can guide ChatGPT’s reasoning (Step 3). To inform this decision, APT-Pipe iteratively tests prompt variations on a small subset of manually annotated data (typically a few hundred samples) to learn the best combination of information to include in the final prompt. Once complete, the prompt can be used on the remaining dataset.
@@ 3.1. Step 1: Preparing Initial Prompt
Figure 1 illustrates the process to prepare the initial prompts. The prompt engineer must first input an initial prompt using a JSON syntax. We use JSON because encoded notations can facilitate language models’ ability to understand and respond with more precise answers (Zhou et al., 2022; Yu et al., 2023a). Moreover, designing prompts as JSONs is convenient for prompt augmentation, as APT-Pipe can easily embed additional information using suitable key-value pairs (Yu et al., 2023a; Kumar et al., 2023). Note, we have also experimented with various commonly applied prompt templates, i.e. natural language (Bang et al., 2023), dictionary (Zhu et al., 2023), cloze question (Hu et al., 2021), and JSON (Zhuo et al., 2023). Based on our findings, we observe that JSON prompts generate more precise responses in a consistent format. Although the engineer could input any preferred JSON template, by default, we rely on the following template:
@@ 3.2. Step 2: Prompt-tuning with Few-shot
Figure 2. Flow chart of Step 2.
@@
APT-Pipe first conducts exemplar selection. To select the samples to include from the manually annotated set, we rely on the method proposed by Yang et al. (Yang et al., 2022). Specifically, we select the text samples from the dataset with the highest textual similarity with the text to classify. Thus, we use OpenAI’s text-embedding-ada-002 (Neelakantan et al., 2022) model to generate the embeddings for each data sample in the dataset and then calculate the pairwise cosine similarity between the embedding of each exemplar and the embedding of the text to classify. Consequently, we focus on the top $n$ exemplars with the highest cosine similarity. Note, $n$ is configurable, allowing the user to select a range of potential values to experiment on. By default, $n=5$ as a five-shot setting is commonly used in related works (Han et al., 2022; Ma et al., 2021). We include the $n$ examples in the template as follows:
@@ 3.3. Step 3: Prompt-tuning with NLP Metrics
Figure 3. Flow chart of Step 3.
@@
Termination Condition:  The loop will continue until all metrics in Step 3.3 have been tested. Thus, the number of iterations is limited by the number of metrics. Importantly, by ranking metrics using XGBoost, we avoid having to experiment with every permutation. Although this cannot be guaranteed to bring optimum performance for all datasets, our experiments have shown that this is an effective estimate. Once all metrics have been tested, APT-Pipe will return a final prompt containing all the chosen metrics as JSON attributes. This becomes the fine-tuned prompt that the prompt engineer will use.
@@ 4. Experimental Settings
In this section, we illustrate the selection of social computing datasets for experiment, evaluation metrics, and our pipeline implementation.
@@ 4.1. Task Domains and Datasets
Our experiments cover twelve public datasets drawn from three significant domains as highlighted by reviews on text classification (Minaee et al., 2021; Küçük and Can, 2020; Crothers et al., 2023):
@@ 4.2. Evaluation Metrics
  • Time cost: The time it takes for ChatGPT to return an annotation response. We normalize the request time by comparing it to the request time of a null prompt sent before. Here, the normalization is done to remove the influence of the network transmission. Afterwords, we divide the normalized request time by the number of words in the response. This metric can be interpreted as the time required for ChatGPT to generate each token.
@@ 4.5. Experimental Implementation
ChatGPT setting:  We have built the full APT-Pipe tool, and integrated it with gpt-3.5-turbo. We request access ChatGPT through OpenAI’s API with parameter temperature set to 00 to make the response focused and offer deterministic. [^1] https://platform.openai.com/docs/api-reference/chat APT-Pipe will be made publicly available on github. [^2] Redacted due to anonymous submission requirements.
@@ 5.2. Effectiveness of ChatGPT’s Responses
As implied in relevant studies, two challenges for applying prompt-tuning for data annotation tasks are that (i) ChatGPT may generate responses in arbitrary formats (Amin et al., 2023; Guo et al., 2023); and (ii) tuning methods could increase the time cost for generating annotations (Shi and Lipani, 2023). We next assess whether APT-Pipe also faces these two difficulties.

### 20 Feb 2024 v4

# APT-Pipe: A Prompt-Tuning Tool for Social Data Annotation using ChatGPT

@@ Abstract.
Recent research has highlighted the potential of LLMs, like ChatGPT, for performing label annotation on social computing text. However, it is already well known that performance hinges on the quality of the input prompts. To address this, there has been a flurry of research into prompt tuning — techniques and guidelines that attempt to improve the quality of prompts. Yet these largely rely on manual effort and prior knowledge of the dataset being annotated. To address this limitation, we propose APT-Pipe, an automated prompt-tuning pipeline. APT-Pipe aims to automatically tune prompts to enhance ChatGPT’s text classification performance on any given dataset. We implement APT-Pipe and test it across twelve distinct text classification datasets. We find that prompts tuned by APT-Pipe help ChatGPT achieve higher weighted F1-score on nine out of twelve experimented datasets, with an improvement of 7.01% on average. We further highlight APT-Pipe’s flexibility as a framework by showing how it can be extended to support additional tuning mechanisms.
Human computation, prompt-tuning, large language models
journalyear: 2024copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589334.3645642isbn: 979-8-4007-0171-9/24/05ccs: Computing methodologies Natural language processing
@@ 1. Introduction
Data annotation commonly relies on human intelligence and suffers from challenges related to annotator quality, data comprehension, domain knowledge, and time cost (Haq et al., 2022; Huang et al., 2016; Veit et al., 2016; Chalkidis et al., 2020). Recent research has highlighted that prompt-based LLM applications, like ChatGPT, have the potential to substitute human intelligence as a data annotator for social computing tasks (Zhu et al., 2023), e.g., sentiment analysis (Bang et al., 2023; Zhu et al., 2023) and hate speech detection (Huang et al., 2023). However, such models are heavily impacted by prompt quality, whereby poorly tuned prompt can severely damage annotation accuracy (Shum et al., 2023). This creates the need for tools that can help researchers design and evaluate effective prompts.
@@
To date, most prompt-tuning for annotation tasks is done either manually (Liao et al., 2022) or by injecting prior knowledge of the dataset (Gilardi et al., 2023; Huang et al., 2023). Prompt-tuning for ChatGPT, however, usually introduces demands on expertise that can be hard to address. Thus, we argue that an automated prompt-tuning tool for ChatGPT tasks could remove a significant manual load related to learning prior knowledge of the dataset to annotate. Researchers have already begun to propose techniques for such automation (Petroni et al., 2019; Min et al., 2021; Bang et al., 2023). Here, we seek to integrate and supplement these approaches within a single pipeline. We focus on three early-stage ideas: introducing standardized templates that (almost) guarantees consistent responses; introducing few-shot examples into prompts to guide annotations; and embedding additional text metadata to augment raw text information. Implementing these simple concepts raises key challenges. First, ChatGPT can generate unpredictable and unstructured responses that make parsing annotations difficult (as its responses lack a standard format). Although cloze-style prompts can improve responses’ parsability, such templates usually introduce a trade-off on ChatGPT’s time efficiency (Petroni et al., 2019). Second, although preliminary evidence suggests that performance can be improved by including few-shot examples of annotation (Min et al., 2021), manual exemplar selection can be time consuming. Thus, techniques need to be developed to automatically identify suitable few-shot examples for the prompt. Third, although embedding natural language processing (NLP) metrics in prompts (Bang et al., 2023) has been shown to increase performance, the selection of such metrics requires domain specific expertise. Thus, techniques must be developed to select the best permutations of NLP metrics to include in prompts. Importantly, the suitability of the above approaches may differ on a per-dataset basis, making a single approach impractical.
@@
To address these challenges, we propose the Automatic Prompt-Tuning Pipeline (APT-Pipe), an extensible prompt-tuning framework for improving text-based annotation task performance. APT-Pipe only requires humans to annotate a small sample subset of data to offer a robust ground truth. It then exploits this annotated subset to automatically tune and test prompts with the goal of improving ChatGPT’s classification performance. APT-Pipe addresses the above three challenges using a three-step prompt-tuning pipeline. It first generates JSON template prompts that can make ChatGPT respond consistently in a desired format. APT-Pipe then automatically computes the most suitable few-shot examples to include in prompts. Finally, through a number of iterations, APT-Pipe tests a variety of prompt configurations with NLP metrics to identify which work best for the particular dataset. These three steps are derived from our experience and state-of-the-literature. However, we acknowledge that many other prompt-tuning techniques are likely to emerge in the coming years. Consequently, we design APT-Pipe in an extensible and modular fashion, allowing other researchers to easily “plug-in” additional tuning strategies.
@@
We evaluate APT-Pipe across twelve text classification datasets drawn from three domains in social computing. Our results shows that prompts tuned by APT-Pipe can improve ChatGPT’s classification performance on nine out of twelve datasets, with a 7.01% increase of weighted F1-score on average. Moreover, prompts tuned by APT-Pipe can decrease ChatGPT’s computing time for classification, while providing over 97% of all responses in a consistent and easy-to-parse format. Finally, we highlight APT-Pipe’s flexibility by extending its framework with two state-of-art tuning mechanisms called Chain-of-Thought and Tree-of-Thought. The contributions of this paper can be summarized as follows:
@@
  • We show that APT-Pipe prompts generate annotations in a consistent format and take less time compared to existing prompt designs. On average, 97.08% responses generated by APT-Pipe are parsable and it decreases the time cost by over 25%.
  • We highlight APT-Pipe’s extensibility by introducing two additional prompt-tuning techniques: Chain of Thought (CoT) and Tree of Thought (ToT). CoT improves APT-Pipe’s F1-scores on five datasets by 4.01%, and ToT improves APT-Pipe’s F1-scores on six datasets by 3.49%. Importantly, we show that APT-Pipe can learn these improvements on a per-dataset basis, automatically tuning the prompts without human intervention.
@@ 2.1. LLMs in Social Computing
Social computing research relies extensively on human-annotated datasets. Such annotations are time consuming and often costly. The release of LLM tools such as ChatGPT (Bang et al., 2023; Guo et al., 2023) has uncovered a range of possibilities to automate text data labeling tasks due to their robust performance in natural language comprehension and reasoning (Kojima et al., 2022). Recently, some studies (Sallam et al., 2023; Huang et al., 2023) have employed LLMs to assist in text annotation in social computing. Zhu et al. (Zhu et al., 2023) evaluate the potential of LLMs to annotate five different types of data. Sallam et al. (Sallam et al., 2023) employ LLMs to discriminate misinformation about vaccines on social media platforms. Huang et al. (Huang et al., 2023) exploit LLMs to annotate implicit hateful tweets and report nearly 80% accuracy. However, work has revealed that the annotation performance of LLMs varies across different LLMs and different domains and datasets (Aiyappa et al., 2023; Shen et al., 2023). The above literature demonstrates that, although it is feasible to employ LLMs for text annotation tasks, the generality and reliability of LLMs need to be improved.
@@ 2.2. Prompt-Tuning for Text Annotation
In existing LLM-based text annotation methods (Alizadeh et al., 2023; Luo et al., 2023; Kuzman et al., 2023; Reiss, 2023), several prompt-tuning techniques have been utilized to enhance the performance of LLMs. Alizadeh et al. (Alizadeh et al., 2023) investigate the performance difference between LLMs and crowd-workers in text annotation tasks. They employ natural language formatted prompts in zero-shot tests. For few-shot tests, they employ the Chain-of-Thought (CoT) (Wei et al., 2022) prompting technique, where LLMs are provided with the question and step-by-step reasoning answers as examples. Kuzman et al. (Kuzman et al., 2023) conduct genre identification tasks using LLMs by feeding prompts with manually defined task descriptions, and manually extracting the answers from responses. Reiss (Reiss, 2023) applies LLMs to text annotation tasks by manually designing 10 prompts, feeding them to LLMs, and evaluating their performance. We observe that most existing prompt-tuning methods require the participation of domain expertise and human judgement, which hinders the utilization of LLMs in automated annotation tasks. Using ChatGPT as a case-study, we focus on proposing a pipeline that can automatically perform prompt tuning.
@@ 3. APT-Pipe Design
We now describe our pipeline for tuning a text classification prompt. The pipeline’s goal is to generate a tuned prompt to improve ChatGPT’s classification performance on a per-dataset basis. The pipeline operates in discrete steps, and we have designed our framework in an extensible fashion, allowing future modifications to be “plugged-in” as new steps. The current implementation has three steps. First, the prompt engineer must personalize a JSON-based prompt following our template (Step 1). Following this, the pipeline searches for suitable few-shot learning examples to include in the prompt (Step 2) and augmentary text metrics that can guide ChatGPT’s reasoning (Step 3). To inform this decision, APT-Pipe iteratively tests prompt variations on a small set of manually annotated data to learn the best combination to include in the final prompt. Once complete, the prompt can be used on the remaining dataset.
@@ 3.1. Step 1: Preparing Initial Prompt
Figure 1 illustrates the process to prepare the initial prompts. The prompt engineer must first input an initial prompt using a JSON syntax. We use JSON because encoded notations can facilitate language models’ ability to understand and respond with more precise answers (Zhou et al., 2022; Yu et al., 2023a). Moreover, designing prompts as JSONs is convenient for prompt augmentation, as APT-Pipe can easily embed additional information using suitable key-value pairs (Yu et al., 2023a; Kumar et al., 2023). Note, we have also experimented with other common prompt templates, i.e. natural language (Bang et al., 2023), dictionary (Zhu et al., 2023), and cloze question (Hu et al., 2021). We observe that JSON prompts generate more precise responses in a consistent format.
Although the engineer could input any preferred JSON template, by default, we rely on the following template:
@@ 3.2. Step 2: Prompt-tuning with Few-shot
Figure 2. Flow chart of Step 2 (prompt-tuning with few-shot learning).
@@
APT-Pipe first conducts exemplar selection. To select the samples to include from the manually annotated set, we rely on the method proposed by Yang et al. (Yang et al., 2022). Specifically, we select the text samples from the dataset with the highest textual similarity with the text to classify. We use OpenAI’s text-embedding-ada-002 (Neelakantan et al., 2022) model to generate the embeddings for each data sample in the dataset and then calculate the pairwise cosine similarity between the embedding of each exemplar and the embedding of the text to classify. Consequently, we focus on the top $n$ exemplars with the highest cosine similarity. Note, $n$ is configurable, allowing the user to select a range of potential values to experiment on. By default, $n=5$, which is a five-shot setting commonly used in related works (Han et al., 2022; Ma et al., 2021). We include the $n$ examples in the template as follows:
@@ 3.3. Step 3: Prompt-tuning with NLP Metrics
Figure 3. Flow chart of Step 3 (prompt-tuning with NLP metrics).
@@
Termination Condition:  The loop will continue until all metrics in Step 3.3 have been tested. Thus, the number of iterations is limited by the number of metrics. By ranking metrics using XGBoost, we avoid having to experiment with every permutation. Although this cannot guarantee to bring optimum performance for all datasets, our experiments have shown that this is an effective estimate. Once all metrics have been tested, APT-Pipe will return a final prompt containing all the chosen metrics as JSON attributes. This becomes the fine-tuned prompt that the prompt engineer will use.
@@ 4. Experimental Settings
In this section, we outline our experimental setup, covering evaluation datasets, metrics, and our pipeline implementation.
@@ 4.1. Task Domains and Datasets
Our experiments cover twelve public datasets drawn from three significant domains, as highlighted by reviews on text classification (Minaee et al., 2021; Küçük and Can, 2020; Crothers et al., 2023):
@@ 4.2. Evaluation Metrics
  • Time cost: The time it takes for ChatGPT to return an annotation response. We normalize the request time by comparing it to the request time of a null prompt sent before. Here, the normalization is done to remove the influence of the network transmission. Afterwords, we divide the normalized request time by the number of tokens in the response. This metric can be interpreted as the time required for ChatGPT to generate each token.
@@ 4.5. Experimental Implementation
ChatGPT setting:  We have built the full APT-Pipe tool, and integrated it with gpt-3.5-turbo. We request access to ChatGPT through OpenAI’s API with parameter temperature set to 00 to make the response focused and offer deterministic. [^1] https://platform.openai.com/docs/api-reference/chat APT-Pipe will be made publicly available on github. [^2] https://github.com/James-ZYM/APT-Pipe
@@ 5.2. Effectiveness of ChatGPT’s Responses
Two challenges for applying prompt-tuning for data annotation tasks are that (i) ChatGPT may generate responses in arbitrary formats (Amin et al., 2023; Guo et al., 2023); and (ii) tuning methods could increase the time cost for generating annotations (Shi and Lipani, 2023). We next assess whether APT-Pipe also faces these two difficulties.